# STAGNet: A Spatio-Temporal Graph and LSTM Framework for Accident Anticipation

### v1

@@ 5 Conclusion
In this work, we demonstrate that state-of-the-art graph neural networks for accident anticipation can be substantially improved by using SlowFast networks to extract global, spatio-temporal features of video frames and aggregating them through an LSTM network. Our experiments on benchmark datasets show that the proposed STAGNet model can anticipate ego-involved accidents earlier and more accurately compared to existing methods. Our approach performs better than others even when tested on a dataset different from the one used to train it (cross-dataset evaluation). However, for all algorithms tested, the cross-dataset average precision and mean time-to-accident are notably lower than those achieved when cross-validated on the same dataset. This suggests that further research is needed to investigate and improve the generalization of these models to new environments and traffic conditions.
We note that our model (and other models compared here) is currently not optimized for execution speed. In future work, we will explore techniques to make it more computationally efficient. We also plan to enhance vehicle localization by moving beyond bounding box–based approaches, which can leave empty spaces and fail to tightly encapsulate vehicles, especially along object edges. Utilizing segmentation techniques would allow for more precise delineation of vehicle boundaries, thereby improving detection accuracy. However, this could lead to a reduction in inference speed. Therefore, the trade-offs need to be considered carefully. Additionally, incorporating depth information or utilizing the cameras’ intrinsic matrix could enable more accurate estimation of the distance between vehicles, facilitating better scene understanding and potentially improving the robustness of accident prediction models.

### v2

@@ 4.2.5 Analysis of Computational Efficiency and Execution Speed
STAGNet-Lite - 0.045 (SlowFast) 0.002 21.317
* Time for saliency/driver-attention computation.
Table 9: FLOP counts of different methods. Counts for different stages are reported in GFLOPs/frame. The counts were obtained using the Python library THOP [THOP].
Method GFLOPs
Object Detection & Feature Extraction Global Feature Extraction Model Inference Total
UString 322.46 15.47 (VGG16) 10.49 348.42
Graph (Graph) 41.75 (I3D) 2.14 366.35
AAT-DA 15.47 (VGG16) 1.05 461.96*
STAGNet 50.58 (SlowFast) 2.20 375.24
STAGNet-Lite - 50.58 (SlowFast) 0.17 50.75
* The total includes the 122.98 GFLOPs used for saliency/driver-attention computation.
@@ 5 Conclusion
In this work, we demonstrate that state-of-the-art graph neural networks for accident anticipation can be substantially improved by using SlowFast networks to extract global, spatio-temporal features of video frames and aggregating them through an LSTM network. Our experiments on benchmark datasets show that the proposed STAGNet model can anticipate ego-involved accidents earlier and more accurately compared to existing methods. Our approach outperforms the others even when tested on a dataset different from the one used for training (cross-dataset evaluation). However, for all algorithms tested, the cross-dataset average precision and mean time-to-accident are notably lower than those achieved when cross-validated on the same dataset. The reasons for this drop in performance are not clear from the experiments we conducted. This suggests that further research is needed to investigate and improve the generalization of these models to new environments and traffic conditions.
We note that our full STAGNet model and the other models compared here are currently not optimized for execution speed. Our pruned model, STAGNet-Lite, operates at a much faster speed while being almost as effective as the full model in detecting ego-involved accidents. In future work, we will further explore techniques to make the models more computationally efficient. We also plan to enhance vehicle localization by moving beyond bounding box–based approaches, which can leave empty spaces and fail to encapsulate vehicles tightly. Utilizing segmentation techniques would enable more precise delineation of vehicle boundaries, thereby improving detection accuracy. However, this could lead to a reduction in inference speed. Therefore, the trade-offs need to be considered carefully. Additionally, incorporating depth information or utilizing the cameras’ intrinsic matrix could enable more accurate estimation of the distance between vehicles, facilitating better scene understanding and potentially improving the robustness of accident prediction models.